0=3x*3x-5x+11
x*x-6x+9=0
5+3*3-16+4
12+20/4-5
the area of a circle is 20(3.14) find the radius
the radius of a circle is 9. find circumference

Answer by Nate(3500)

AREA = (pi)r^2
20(3.14) = (3.14)r^2
20 = r^2
sqrt(20) = r
~~~~
r = 9
Circumference = 2(3.14)*9 = 56.52